# Krylov fractality and complexity in generic random matrix ensembles

### v1

@@ Abstract
Krylov space methods provide an efficient framework for analyzing the dynamical aspects of quantum systems, with tridiagonal matrices playing a key role. Despite their importance, the behavior of such matrices from chaotic to integrable states, transitioning through an intermediate phase, remains unexplored. We aim to fill this gap by considering the properties of the tridiagonal matrix elements and the associated basis vectors for appropriate random matrix ensembles. We utilize the Rosenzweig-Porter model as our primary example, which hosts a fractal regime in addition to the ergodic and localized phases. We discuss the characteristics of the matrix elements and basis vectors across the three (ergodic, fractal, and localized) regimes and introduce tools to identify the transition points. The exact expressions of the Lanczos coefficients are provided in terms of $q$-logarithmic function across the full parameter regime. The numerical results are corroborated with analytical reasoning for certain features of the Krylov spectra. Additionally, we investigate the Krylov state complexity within these regimes, showcasing the efficacy of our methods in pinpointing these transitions.
@@
This manuscript addresses phase transitions within a prototypical random matrix ensemble, the Rosenzweig-Porter model (RP) Rosenzweig and Porter (1960, 1960); Altland et al. (1997a); Khaymovich and Kravtsov (2021); Pino et al. (2019), employing Krylov space techniques Parker et al. (2019); Balasubramanian et al. (2022); Nandy et al. (2024). These techniques have gained prominence due to their simplicity and wide applicability; see Barbón et al. (2019); Dymarsky and Gorsky (2020); Rabinovici et al. (2021, 2022a); Dymarsky and Smolkin (2021); Caputa et al. (2022); Bhattacharjee et al. (2022a); Hörnedal et al. (2022); Bhattacharya et al. (2022); Rabinovici et al. (2022b); Bhattacharjee et al. (2022b); Liu et al. (2023); Bhattacharjee et al. (2023); Avdoshkin et al. (2024); Camargo et al. (2023); Erdmenger et al. (2023) for the initial explorations and Nandy et al. (2024) for a detailed review and the references therein. By transforming the Hamiltonian into a tridiagonal matrix, one analyzes its elements, known as Lanczos coefficients, which comprise the Krylov spectrum Parker et al. (2019); Balasubramanian et al. (2022). Such techniques have been employed in a few simplified cases across different phases Scialchi et al. (2024); Alishahiha and Vasli (2024); Menzler and Jha (2024); Cohen et al. (2024). The RP model is particularly intriguing as it exhibits a spectral fractal phase, bridging the gap between ergodic and localized phases. This complex phase structure offers a unique opportunity to apply Krylov space methodologies. Considering the statistics of the Lanczos coefficients and the degree of localization of the Krylov vectors, we propose approximate Ansätze that adeptly characterizes the behavior of the Lanczos coefficients during the transitional phases. Our analytical discourse substantiates the observed behaviors. We analyze the inverse participation ratio (IPR) Evers and Mirlin (2008) of Krylov vectors to assess the degree of localization, uncovering the IPR’s ability to differentiate among the three distinct phases. We also investigate the complexity associated with the spread of a Gibbs state through these phases, identifying specific features that sharply delineate phase boundaries.
Rosenzweig-Porter model: The Rosenzweig-Porter (RP) model Rosenzweig and Porter (1960) is given by the $N\times N$ Hamiltonian of the form Venturelli et al. (2023)
@@
where $A$ is a diagonal matrix, drawn from a normal distribution with zero mean and unit variance, i.e., $\braket{a_{ii}}^{2}=1$, and $B$ is a matrix from Gaussian Orthogonal Ensemble (GOE), drawn from a normal distribution with zero mean (for both diagonal and off-diagonal elements) and variance $\braket{b_{ii}}^{2}=1$ and $\braket{b_{ij}}^{2}=1/2$ for $i\neq j$, respectively. Here $\gamma\geq 0$ is a parameter that effectively controls the dominant behavior of $A$ and $B$. It gives rise to the corresponding delocalized or ergodic ($\gamma\in[0,1]$), fractal ($\gamma\in[1,2]$), and localized ($\gamma\in[2,\infty)$) phases, which leave imprints on the eigenspectrum of $H$. In the large $N$ limit, one can obtain the $\braket{r}$ value statistics of the mid-spectrum Atas et al. (2013). Specifically, one obtains the Wigner-Dyson distribution for $\gamma=[0,2)$ while the Poissonian distribution appears in the $\gamma=(2,\infty)$ limit Altland et al. (1997b); see Supplemental Material (SM1) sup for further details. On the other hand, the fractal phase is characterized by the fractal dimension $\mathcal{D}$, defined through the scaling of the Inverse Participation Ratio (IPR) as $\mathrm{IPR}=\sum_{n}|\braket{n}{\psi}|^{4}\sim N^{-\mathcal{D}}$ Buijsman and Lev (2022). Here |$\ket{\psi}$ is a state in the eigenbasis and |$\ket{n}$ is a computational basis element, and the overlap captures the spread of the eigenstate in the computational basis. The ergodic and localized regime correspond to constant fractal dimensions, i.e., $\mathcal{D}=1$ and $\mathcal{D}=0$, respectively, while in the fractal phase, the fractal dimension linearly decreases as $\mathcal{D}=2-\gamma$ Kravtsov et al. (2015).
@@
Here, the sets $\{\mathsf{a}_{n},\mathsf{b}_{n}\}$ are referred to as the Lanczos coefficients, and the set {|$\{\ket{K_{n}}\}$ constitutes the Krylov basis, which together forms the Krylov spectrum Balasubramanian et al. (2022); see Nandy et al. (2024) for a detailed review. For numerical purposes, it is often useful to consider the Hessenberg decomposition Golub et al. (1979), which uses the Householder reflections Householder (1958) instead of the Lanczos algorithm. This produces identical results for Lanczos coefficients, subject to some redefinition of Krylov basis vectors. Curiously, the density of the states can be well approximated from the statistics of the Lanczos coefficients $\mathsf{a}(x)$ and $\mathsf{b}(x)$ where $x=n/N$ in the large $N$ limit as Kuijlaars and Van Assche (1999); Balasubramanian et al. (2023a, b)
@@
Figure 1: (Left) The Lanczos coefficients for the RP model with $N=4096$ (and $100$ averaging) as a function of $x=n/N$, where $n$ is the index of the Lanczos coefficient. The black dashed lines are the fitted results given by (6) (top) and (8) (bottom) at the two extremes. (Right) The behavior of the fitting parameters $u$ and $v$ as functions of $\gamma$ and $N$.
Statistics of Lanczos coefficients: In a manner akin to the conventional eigenspectrum analysis, we focus on the behavior of the Lanczos coefficients for the Hamiltonian (1) throughout all phases, including the critical points of fractal and localization transitions. Figure 1 shows the behavior of the Lanczos coefficients in the RP model as a function of $\gamma$. As is evident from this, there is a transition in the behavior of the Lanczos coefficients as $\gamma$ transits from the ergodic to the localized phase. The behavior in the ergodic phase (where the matrix is full GOE) is known analytically Zuker et al. (2001); Balasubramanian et al. (2023a) and given by
@@
The symmetry of the energy in the density of states requires $\mathsf{a}(x)=0$ upon the ensemble average of the Hamiltonian. In the localized phase $\gamma\gtrsim 2$, the nature of the function is previously unknown. In this paper, we ascertain that such a form is given by
@@
where $d=\log_{2}N$ and $N$ being the matrix dimension. Here “nib” stands for the inverse of the shifted binomial function $\text{bin}(x,d)=2^{-d}\binom{d}{d(1/2-x)}$ Zuker et al. (2001). Such structure has been previously studied in the context of nuclear spectra Molina et al. (2005); Zuker et al. (2001), yet remains largely unknown. The inverse binomial is ill-defined at the origin, while away from $x\sim 0$, it can be approximated to a simpler form
@@
with corresponding Gaussian density of states of zero mean and variance $\xi^{2}$, obtained through (3); see SM2 sup . We numerically adjust the factor $\xi$ to make the analytical expression (8) coincide with the value of $x$ where the numerical $\mathsf{b}(x)$ is maximum. We find excellent agreement with the numerical results with $\xi\simeq 1/2$, shown by the black dashed line (green region) in Fig. 1 (left).
@@
given by the $q$-logarithm, introduced by Tsallis Tsallis (1994). It reduces to the usual logarithm for $q\rightarrow 1$, i.e., in the Poisson limit. Here, $p\equiv p(\gamma,N)$ and $q\equiv q(\gamma,N)$ weakly depend on the system size $N$. We scale the coefficients appropriately to lie between zero and unity. Importantly, (9) reduces to (6) and (8) at $q\rightarrow 0$ and $q\rightarrow 1$ limits, which are proxy for the $\gamma\rightarrow 0$ and $\gamma\gtrsim 2$ limits, respectively. The corresponding limits of $p$ are $p\simeq 1$ and $p\simeq\xi^{2}/2$. We also note that the Ansatz (9) is not unique, and an alternate Ansatz is discussed in the SM3 sup .
@@
Curiously, the density of states corresponding to (9) can be obtained by solving the integral equation (3). It is given by the following form
@@
Behavior in deep localized regime - analytic arguments: A notable aspect of the RP model in its tridiagonal form is the persistence of non-zero off-diagonal elements even in the deeply localized regime, as evidenced by Fig. 1. At first glance, this seems counterintuitive, especially when considering that at high values of $\gamma$, the primary contribution should theoretically arise from the diagonal matrix $A$. Yet, our numerical findings indicate that the off-diagonal elements remain finite and are of comparable magnitude to those found in the ergodic regime, a phenomenon we explore below.
We first examine the response of an orthogonal transformation to the RP model. For a Hamiltonian characterized by the distribution with $\langle H_{ij}\rangle=0$ and $\langle H^{2}_{ij}\rangle=\alpha\delta_{ij}+\beta(1-\delta_{ij})$, an orthogonal transformation via a matrix $C$ keeps the expectation value unchanged $\langle\tilde{H}_{ij}\rangle=0$ while transforming the variance to $\langle\tilde{H}^{2}_{ij}\rangle=\beta(1+\delta_{ij})+(\alpha-2\beta)\sum_{l}c^{2}_{li}c^{2}_{lj}$; see SM5 sup for details. Here $\tilde{H}$ denotes the transformed Hamiltonian, and $c_{ij}$ represents the elements of $C$. Although this matrix can be arbitrary, we limit it to a special orthogonal matrix $M$ that executes the Householder transformation on $H$. The elements of $M$ can be approximated using the distribution of $H_{ij}$ sup . The variance of the distribution of the elements of the matrix $\tilde{H}$ after a single orthogonal transformation is given by
@@
Krylov IPR: To gain further insight, we address the extension of localization in Krylov space by computing the Inverse Participation Ratio (IPR) for the Krylov subspace. Instead of taking the usual eigenvectors, this amounts to considering the Krylov vectors corresponding to an initial localized state. It is widely understood that for quantum mechanical systems, Krylov space localization agrees with real space localization. The eigenvector IPR has been extensively studied in the RP model Cadez et al. (2024). We extend this study to the Krylov IPR and determine the appropriate scaling that allows us to distinguish between the three phases of the RP model. We define the Krylov IPR as the overlap between the computational basis state and the Krylov state
@@
where $|\varphi_{k}\rangle$ stands for the $k^{\text{th}}$ concerned Krylov vector and |$\ket{n}$ is a computational basis element. The logarithm of the $\ell=1$ case is often known as the participation entropy.
The Krylov state |$\ket{\varphi_{k}}$ can be expressed in terms of the Hamiltonian eigenstates as |$\ket{\varphi_{k}}=\sum_{m=1}^{N}\eta^{k}_{m}\ket{\psi_{m}}$, where |$\ket{\psi_{m}}$ is an eigenstate of the Hamiltonian $H$. The coefficients $\eta^{k}_{m}$ satisfy a three-term recursion relation; see SM6 sup . Thus, $\mathrm{IPR}^{\ell}_{\text{K}}$ can be rewritten as
$\mathrm{IPR}^{\ell}_{\text{K}}(\varphi_{k})=\sum_{n=1}^{N}\Big{|}\sum_{m=1}^{N}\eta^{k}_{m}s^{m}_{n}\Big{|}^{2\ell}\,,$ (13)
where $s^{m}_{n}=\braket{n}{\psi_{m}}$. The usual IPR for the eigenstate $|\psi_{k}\rangle$ is then obtained by simply setting $\eta_{m}^{k}=1$.
The usual eigenstate IPR information is encoded in the Krylov IPR. In SM6 sup , the overlap of Krylov vectors and eigenstates is also investigated. It suggests that this quantity can be used as an effective probe for localization and multifractality Khaymovich and Kravtsov (2021). Physically, this is a combined measure of the spread of eigenstates on the computational basis and the spread of the Krylov vectors in the eigenbasis. Note that our definition of Krylov IPR is static and reflects the localization with respect to the $\gamma$ in comparison to the dynamic IPR defined in Bento et al. (2024).
The detailed behavior of Krylov IPR is discussed in the SM6 sup . Here, we are interested in the scaling of the IPR. The scaling is indeed much more apparent for $k=N-1$ as compared to $k=N/2$, i.e., the middle of the Krylov spectrum. The results are presented in Fig. 2, and we compare them with the $\ell=1$ case in the SM6 sup . The scaling exponents are obtained by fitting the linear decrease $\mathrm{IRP}^{2}_{\text{K}}(\varphi_{N-1})\sim N^{-\mathcal{D}_{2}}$ in the fractal region. The following features are of the scaling exponent $\mathcal{D}_{2}$
@@
Up to numerical and finite-size effects, it is apparent that $\mathcal{D}_{2}(\gamma)=\text{constant}>0$ for $\gamma\leq 1$ (i.e., in the metallic phase). In the fractal phase, $\mathcal{D}_{2}(\gamma)\sim-\alpha\gamma+\chi>0$ (where $\alpha$ and $\chi$ are some positive constants) for $1<\gamma\leq 2$ (i.e., in the fractal phase). Finally, we find that $\mathcal{D}_{2}(\gamma)\sim 0$ for $\gamma>2$ (i.e., in the localized phase), indicating strong localization.
The underlying reason behind the strong sensitivity of $\varphi_{N-1}$ (or any $\varphi_{k}$ close to $k=N$) to $\gamma$ comes from the Lanczos algorithm. Since $\varphi_{k}$ is generated by the application of $H^{k}$ on the initial state followed by the subtraction of the contribution of $\varphi_{k-1}$ and $\varphi_{k-2}$, the effect of the full Lanczos spectrum is encoded in the Krylov vectors towards the end of the spectrum. For example, by considering the overlap of all Krylov vectors with the mid-spectrum eigenstate (63), one finds that the larger the value of $k$, the more information about the Lanczos spectrum is encoded in this overlap coefficient. Since we have seen that the full Lanczos spectrum is sensitive to the transition, it is desirable to utilize a Krylov vector that encodes the full spectrum.
Figure 2: (Left) Scaling of the $\mathrm{IPR}^{2}_{\text{K}}$ for $\varphi_{N-1}$ for different values of $N$ as a function of $\gamma$. The scaling with system size is nearly unaffected by $\gamma$. (Right) Scaling exponent $\mathcal{D}_{2}$ of $\mathrm{IPR}^{2}_{\text{K}}$ for $\varphi_{N-1}$ as a function of $\gamma$. In both cases, the exponent is nearly constant in the ergodic regime and decreases linearly in the fractal regime to settle at zero in the localized regime.
@@
Figure 3: (Top) Spread complexity for the RP model (1) in different regimes, initialized by a TFD state at infinite temperature. (Bottom, left) The variation of the peak value of the spread complexity, and (bottom, right) the peak time in different phases of the Hamiltonian (1), initialized by a TFD state at infinite temperature. The system size is $N=500$, variance $\sigma^{2}=1/N$ with $800$ Hamiltonian realizations.
@@
Conclusion and outlook: In this study, we investigate the Rosenzweig-Porter (RP) model, a random matrix model that exhibits extended, fractal, and localized phases. Our exploration utilizes both static and dynamic measures within the Krylov subspace to shed light on the phase transitions among these three distinct states. We present a semi-analytic expression for the Lanczos coefficients that interpolates between the three phases. We introduce the Krylov Inverse Participation Ratio (IPR) and a Krylov fractal dimension, which accurately indicate the two critical transition points. Notably, the edge-spectrum Krylov vectors exhibit stronger sensitivity to fractality, in contrast to those near the center of the spectrum. Dynamic metrics, such as the spread complexity of Krylov states, also reflect these transitions.
@@
Our findings affirm that both static and dynamic Krylov subspace probes are attuned to the transitions from ergodicity through fractality to localization. These probes present robust alternatives to eigenspectrum analysis. This work bridges a significant gap in the literature on Krylov techniques, especially by providing an analytic expression of Lanczos coefficients across the full parameter regime, thereby establishing a standard for detecting transitions from chaotic to non-chaotic states in physical systems via Krylov methodologies.

### v2

@@ Abstract
Krylov space methods provide an efficient framework for analyzing the static and dynamical aspects of quantum systems, with tridiagonal matrices playing a key role. Despite their importance, the behavior of such matrices from chaotic to integrable states, transitioning through an intermediate phase, remains unexplored. We aim to fill this gap by considering the properties of the tridiagonal matrix elements and the associated basis vectors for appropriate random matrix ensembles. We utilize the Rosenzweig-Porter model as our primary example, which hosts a fractal phase in addition to the ergodic and localized regimes. We discuss the characteristics of the matrix elements and basis vectors across these three (ergodic, fractal, and localized) regimes and introduce Krylov space tools to identify the transition points. The exact expressions of the Lanczos coefficients of such tridiagonal matrices are provided in terms of $q$-logarithmic function across the full parameter regime. The numerical results are corroborated with analytical reasoning for certain features of the Krylov spectra. Additionally, we investigate the Krylov state (spread) complexity within these regimes, showcasing the efficacy of our methods in pinpointing these transitions.
@@
This Letter addresses phase transitions within a prototypical random matrix ensemble, employing Krylov space techniques Parker et al. (2019); Balasubramanian et al. (2022); Nandy et al. (2024). These techniques have gained prominence due to their simplicity and wide applicability; see Barbón et al. (2019); Dymarsky and Gorsky (2020); Rabinovici et al. (2021, 2022); Dymarsky and Smolkin (2021); Caputa et al. (2022); Bhattacharjee et al. (2022a); Hörnedal et al. (2022); Bhattacharya et al. (2022); Liu et al. (2023); Bhattacharjee et al. (2022b, 2023); Avdoshkin et al. (2024); Camargo et al. (2023); Erdmenger et al. (2023) for the initial explorations and Nandy et al. (2024) for a detailed review and the references therein. By transforming the Hamiltonian into a tridiagonal form, one analyzes its elements, known as Lanczos coefficients, which comprise the Krylov spectrum Parker et al. (2019); Balasubramanian et al. (2022). Such techniques have been employed in a few particular cases in different phases Scialchi et al. (2024); Alishahiha and Vasli (2024); Menzler and Jha (2024); Cohen et al. (2024). However, the behavior across the transition regime remains unexplored. We address such transition through Krylov space tools in the Rosenzweig-Porter model Rosenzweig and Porter (1960, 1960); Altland et al. (1997); Khaymovich and Kravtsov (2021); Pino et al. (2019) which is particularly intriguing as it exhibits a spectral fractal phase, bridging the gap between ergodic and localized regimes. We propose approximate Ansätze that characterize the statistical behavior of the Lanczos coefficients during the transitional phases. The static inverse participation ratio (IPR) Evers and Mirlin (2008) of Krylov vectors is introduced to assess the degree of localization, uncovering the IPR’s ability to differentiate among the three distinct phases. We also investigate the complexity associated with the spread of a Gibbs state through these phases, identifying specific features that sharply delineate phase boundaries.
Rosenzweig-Porter model: The Rosenzweig-Porter (RP) model Rosenzweig and Porter (1960) is given by the $N\times N$ Hamiltonian Venturelli et al. (2023)
@@
where $A$ is a diagonal matrix, drawn from a normal distribution with zero mean and unit variance, i.e., $\braket{a_{ii}}^{2}=1$, and $B$ is a matrix from the Gaussian orthogonal ensemble (GOE), drawn from a normal distribution with zero mean (for both diagonal and off-diagonal elements) and variance $\braket{b_{ii}}^{2}=1$ and $\braket{b_{ij}}^{2}=1/2$ for $i\neq j$, respectively. Here $\gamma\geq 0$ is a parameter that effectively controls the dominant behavior of $A$ and $B$. It gives rise to the corresponding delocalized or ergodic ($\gamma\in[0,1]$), fractal ($\gamma\in[1,2]$), and localized ($\gamma\in[2,\infty)$) phases, which leave imprints on the eigenspectrum of $H$. In the large-$N$ limit, one can obtain the $\braket{r}$-value statistics of the midspectrum Atas et al. (2013). Specifically, one obtains the Wigner-Dyson distribution for $\gamma=[0,2)$ while the Gaussian distribution appears in the $\gamma=(2,\infty)$ limit Altland et al. (1997); see Supplemental Material (SM1) sup for further details. On the other hand, the fractal phase is characterized by the fractal dimension $\mathcal{D}$, defined through the scaling of the IPR as $\mathrm{IPR}=\sum_{n}|\braket{n}{\psi}|^{4}\sim N^{-\mathcal{D}}$ Buijsman and Lev (2022). Here |$\ket{\psi}$ is a state in the eigenbasis and |$\ket{n}$ is a computational basis element, and the overlap captures the spread of the eigenstate in the computational basis. The ergodic and localized regime correspond to constant fractal dimensions, i.e., $\mathcal{D}=1$ and $\mathcal{D}=0$, respectively, while in the fractal phase, the fractal dimension linearly decreases as $\mathcal{D}=2-\gamma$ Kravtsov et al. (2015).
@@
Here, the sets $\{\mathsf{a}_{n},\mathsf{b}_{n}\}$ are referred to as the Lanczos coefficients, and the set {|$\{\ket{K_{n}}\}$ constitutes the Krylov basis, which together forms the Krylov spectrum Balasubramanian et al. (2022); see Nandy et al. (2024) for a detailed review. For numerical purposes, it is often useful to consider the Hessenberg decomposition Golub et al. (1979), which uses the Householder reflections Householder (1958) instead of the Lanczos algorithm. This produces identical results for Lanczos coefficients, subject to some redefinition of Krylov basis vectors. Curiously, the density of the states (DOS) can be well approximated from the statistics of the Lanczos coefficients $\mathsf{a}(x)$ and $\mathsf{b}(x)$ where $x=n/N$ in the large-$N$ limit as Balasubramanian et al. (2023a, b)
@@
Figure 1: (Left) The Lanczos coefficients for the RP model with $N=4096$ (with $100$ ensemble averages) as a function of $x=n/N$, where $n$ is the index of the Lanczos coefficient. The black dashed lines are the fitted results given by (6) (top) and (8) (bottom) at the two extremes. (Right) The behavior of the fitting parameters $u$ and $v$ as functions of $\gamma$ and $N$.
Statistics of Lanczos coefficients: In a manner akin to the conventional eigenspectrum analysis, we focus on the behavior of the Lanczos coefficients for the Hamiltonian (1) throughout all phases, including the critical points of fractal and localization transitions. We choose the initial state as $(1,0,0,\cdots,0)^{\intercal}$. However, the choice of the initial state does not affect the averaged or mean of the Lanczos spectrum in the bulk, given a sufficient number of ensemble averages of the Hamiltonian is taken. This is because the averaged bulk Lanczos coefficients depend solely on the density of states of the Hamiltonian. However, even after a large number of ensemble averages, the edge coefficients, i.e., $n\sim o(1)$ coefficients, may exhibit a slight dependence on the choice of the initial state in small system sizes. This effect diminishes as the system size increases.
Figure 1 shows the behavior of the rescaled Lanczos coefficients in the RP model as a function of $\gamma$. As the largest eigenvalue scales with system size $N$, selecting larger system sizes will increase the largest eigenvalue, making the Lanczos coefficients dependent on $N$. For example, in the case of GOE, $b_{n}\sim\sqrt{N/2}$ with $n\sim o(1)$ if the variance of GOE matrices is taken as unity. Conversely, if the variances are taken as $2/N$, then $b_{n}\sim 1$ when $n\sim o(1)$, and does not scale with system size. To ensure this fact, we choose appropriate variances for the random matrices so that the Lanczos coefficients are expressed in a system size-independent manner.
We observe a transition in the behavior of the Lanczos coefficients as $\gamma$ transits from the ergodic to the localized phase. The behavior in the ergodic phase (where the matrix is full GOE) is known analytically Zuker et al. (2001); Balasubramanian et al. (2023a) and given by
@@
The symmetry of the energy in the DOS requires $\mathsf{a}(x)=0$ upon the ensemble average of the Hamiltonian. In the localized phase $\gamma\gtrsim 2$, the nature of the function is previously unknown. In this Letter, we ascertain that such a form is given by
@@
where $d=\log_{2}N$ and $N$ is the matrix dimension. Here “nib” stands for the inverse of the shifted binomial function $\text{bin}(x,d)=2^{-d}\binom{d}{d(1/2-x)}$ Zuker et al. (2001). Such structure has been previously studied in the context of nuclear spectra Molina et al. (2005); Zuker et al. (2001), yet remains largely unknown. The inverse binomial is ill defined at the origin, while away from $x\sim 0$, it can be approximated to a simpler form
@@
with corresponding Gaussian DOS of zero mean and variance $\xi^{2}$, obtained through (3); see SM2 sup . Note that the expressions for the Lanczos coefficients are only valid in bulk and the edge corrections in $n\sim O(1)$ should be determined from the moments of the DOS Balasubramanian et al. (2023b). We numerically adjust the factor $\xi$ to make the analytical expression (8) coincide with the value of $x$ where the numerical $\mathsf{b}(x)$ is maximum. We find excellent agreement with the numerical results with $\xi\simeq 1/2$, shown by the black dashed line (green region) in Fig. 1 (left).
@@
given by the $q$ logarithm, introduced by Tsallis Tsallis (1994). It reduces to the usual logarithm for $q\rightarrow 1$, i.e., in the Poisson limit. Here, $p\equiv p(\gamma,N)$ and $q\equiv q(\gamma,N)$ weakly depend on the system size $N$. We scale the coefficients appropriately to lie between zero and unity. Importantly, (9) reduces to (6) and (8) at the $q\rightarrow 0$ and $q\rightarrow 1$ limits, which are proxy for the $\gamma\rightarrow 0$ and $\gamma\gtrsim 2$ limits, respectively. The corresponding limits of $p$ are $p\simeq 1$ and $p\simeq\xi^{2}/2$. Note that the Ansatz (9) is not unique, and an alternate Ansatz is discussed in the SM3 sup .
@@
Curiously, the DOS corresponding to (9) can be obtained by solving the integral equation (3), which is valid in the bulk of the spectrum. We find
@@
Behavior in the deep localized regime - analytic arguments: A notable aspect of the RP model in its tridiagonal form is the persistence of nonzero off-diagonal elements even in the deeply localized regime, as evidenced by Fig. 1. At first glance, this seems counterintuitive, especially when considering that at high values of $\gamma$, the primary contribution should theoretically arise from the diagonal matrix $A$. Yet, our numerical findings indicate that the off-diagonal elements remain finite and are of comparable magnitude to those found in the ergodic regime, a phenomenon we explore below.
We first examine the response of an orthogonal transformation to the RP model. For a Hamiltonian characterized by the distribution with $\langle H_{ij}\rangle=0$ and $\langle H^{2}_{ij}\rangle=\alpha\delta_{ij}+\beta(1-\delta_{ij})$, an orthogonal transformation via a matrix $C$ keeps the expectation value unchanged $\langle\tilde{H}_{ij}\rangle=0$ while transforming the variance to $\langle\tilde{H}^{2}_{ij}\rangle=\beta(1+\delta_{ij})+(\alpha-2\beta)\sum_{l}c^{2}_{li}c^{2}_{lj}$; see SM6 sup for details. Here $\tilde{H}$ denotes the transformed Hamiltonian, and $c_{ij}$ represents the elements of $C$. Although this matrix can be arbitrary, we limit it to a special orthogonal matrix $M$ that executes the Householder transformation on $H$. The elements of $M$ can be approximated using the distribution of $H_{ij}$ sup . The variance of the distribution of the elements of the matrix $\tilde{H}$ after a single orthogonal transformation is given by
@@
Krylov IPR: To gain further insight, we address the extension of localization in Krylov space by computing the inverse participation ratio (IPR) for the Krylov subspace. Instead of taking the usual eigenvectors, this amounts to considering the Krylov vectors corresponding to an initial localized state. It is widely understood that for quantum mechanical systems, Krylov space localization agrees with real space localization. The eigenvector IPR has been extensively studied in the RP model Čadež et al. (2024). We extend this study to the Krylov IPR and determine the appropriate scaling that allows us to distinguish between the three phases of the RP model. We define the Krylov IPR as the overlap between the computational basis state and the Krylov state
@@
where $|\varphi_{k}\rangle$ stands for the $k^{\text{th}}$ Krylov vector and |$\ket{n}$ is a computational basis element. The logarithm of the $\ell=1$ case is often known as the participation entropy.
The Krylov state |$\ket{\varphi_{k}}$ can be expressed in terms of the Hamiltonian eigenstates |$\ket{\psi_{m}}$ as |$\ket{\varphi_{k}}=\sum_{m=1}^{N}\eta^{k}_{m}\ket{\psi_{m}}$. The coefficients $\eta^{k}_{m}$ satisfy a three-term recursion relation; see SM7 sup . Thus, $\mathrm{IPR}^{\ell}_{\text{K}}$ can be further rewritten as $\mathrm{IPR}^{\ell}_{\text{K}}(\varphi_{k})=\sum_{n=1}^{N}\Big{|}\sum_{m=1}^{N}\eta^{k}_{m}s^{m}_{n}\Big{|}^{2\ell}$ where $s^{m}_{n}=\braket{n}{\psi_{m}}$. The usual IPR for the eigenstate $|\psi_{k}\rangle$ is then obtained by simply setting $\eta_{m}^{k}=\delta^{m}_{k}$ i.e., the usual eigenstate IPR information is encoded in the Krylov IPR. In SM7 sup , the overlap of Krylov vectors and eigenstates is also investigated. This indicates that the Krylov IPR can serve as a robust indicator for both localization and multifractality Khaymovich and Kravtsov (2021). Physically, this corresponds to the spread of eigenstates on the computational basis and the spread of the Krylov vectors in the eigenbasis. Note that our definition of the Krylov IPR is static and reflects the localization with respect to the $\gamma$ in comparison to the dynamic IPR defined in Bento et al. (2024).
The detailed behavior of the Krylov IPR is discussed in the SM7 sup . Here, we are interested in the scaling of the IPR, which is much more apparent for $k=N-1$ than $k=N/2$, i.e., the middle of the Krylov spectrum. The results are presented in Fig. 2, and we compare them with the $\ell=1$ case in SM7 sup . The IPR is higher in the localized regime compared to the ergodic and fractal phases. The scaling exponent is obtained by fitting the linear decrease $\mathrm{IRP}^{2}_{\text{K}}(\varphi_{N-1})\sim N^{-\mathcal{D}_{2}}$ in the fractal region. We observe the following features of the scaling exponent
@@
Up to the numerical and the finite-size effects, $\mathcal{D}_{2}(\gamma)=\text{constant}>0$ for $\gamma\leq 1$ (ergodic phase). In the fractal phase, it linearly decreases $\mathcal{D}_{2}(\gamma)\sim-\alpha\gamma+\chi>0$ (where $\alpha$ and $\chi$ are some positive constants) for $1<\gamma\leq 2$ (fractal phase) while $\mathcal{D}_{2}(\gamma)\sim 0$ for $\gamma>2$ (localized phase), indicating strong localization.
The underlying reason behind the strong sensitivity of $\varphi_{N-1}$ (or any $\varphi_{k}$ close to $k=N$) to $\gamma$ comes from the Lanczos algorithm. In the context of matrix diagonalization, the Lanczos algorithm converges from both extreme ends, making it particularly effective for finding the smallest or largest few eigenvalues and eigenvectors. However, to obtain the eigenvalues and eigenvectors in the middle of the spectrum, the algorithm must effectively span the entire basis. Since $\varphi_{k}$ is generated by the application of $H^{k}$ on the initial state followed by the subtraction of the contribution of $\varphi_{k-1}$ and $\varphi_{k-2}$, the effect of the full Lanczos spectrum is encoded in the Krylov vectors toward the end of the spectrum. For example, by considering the overlap of all Krylov vectors with the mid-spectrum eigenstate, one finds that the larger the value of $k$, the more information about the Lanczos spectrum is encoded in this overlap coefficient. Since we have seen that the full Lanczos spectrum is sensitive to the transition, it is desirable to utilize a Krylov vector that encodes the full spectrum. This also explains why $\varphi_{N-1}$ is highly sensitive to mid-spectrum eigenstates.
Figure 2: Left: Scaling of the $\mathrm{IPR}^{2}_{\text{K}}$ for $\varphi_{N-1}$ for different values of $N$ as a function of $\gamma$. The scaling with system size is nearly unaffected by $\gamma$. The numerical instability at low $\gamma$ arises due to the finite size of the system, which reduces with increasing system size. Right: Scaling exponent $\mathcal{D}_{2}$ of $\mathrm{IPR}^{2}_{\text{K}}$ for $\varphi_{N-1}$ as a function of $\gamma$. In both cases, the exponent is nearly constant in the ergodic regime and decreases linearly in the fractal regime to settle at zero in the localized regime.
@@
Figure 3: Top: Spread complexity for the RP model (1) in different regimes, initialized by a TFD state at infinite temperature. Bottom, left: The variation of the peak value of the spread complexity, and (bottom right) the peak time in different phases of the Hamiltonian (1), initialized by a TFD state at infinite temperature. The system size is $N=500$, variance $\sigma^{2}=1/N$ with $800$ Hamiltonian realizations.
@@
Conclusion and outlook: In this study, we investigate the Rosenzweig-Porter (RP) model, a random matrix model that exhibits extended, fractal, and localized phases. Our exploration utilizes static and dynamic measures within the Krylov subspace to illuminate the phase transitions among these three distinct states. We present a semianalytic expression for the Lanczos coefficients that interpolates between the three phases. We introduce the Krylov inverse participation ratio (IPR) and a Krylov fractal dimension, which accurately indicate the two critical transition points. Notably, the edge-spectrum Krylov vectors exhibit stronger sensitivity to fractality, in contrast to those near the center of the spectrum. Dynamic metrics, such as the spread complexity of Krylov states, also reflect these transitions.
@@
Our findings affirm that both static and dynamic Krylov subspace probes are attuned to the transitions from ergodicity through fractality to localization. These probes present robust alternatives to eigenspectrum analysis. This work bridges a significant gap in the literature on Krylov techniques, especially by providing an analytic expression of Lanczos coefficients across the full parameter regime, thereby establishing a standard for detecting transitions from chaotic to nonchaotic states in physical systems via Krylov methodologies Nandy et al. (2024).